jagr2713 (jagr2713):

Using a directrix of y = -3 and a focus of (2, 1), what quadratic function is created?

jagr2713 (jagr2713):

f(x) = one eighth (x - 2)2 - 1 f(x) = -one eighth (x + 2)2 - 1 f(x) = one half (x - 2)2 + 1 f(x) = -one half (x + 2)2 - 1

jagr2713 (jagr2713):

@freckles @Hero @wio @Zarkon

jagr2713 (jagr2713):

@mathstudent55